Line 1: -3x-3y=6
The slope of the line is
Line 2: 4x+4y=8
The slope of second Line is
What can you infer about the two lines?

A. They are parallel
B.They are perpendicular
C. They are equal
D. None of the above
I just need to check answer to see if it’s right

Answer by Boreal(15235)

put in slope intercept form and look at the slopes
The first is 3y=-3x-6, y=-x-2 so the slope is -1
the second is 4y=-4x+8 y=-x+2 so that slope is -1
They are parallel.